OpenStudy (anonymous):

What is the solution to the equation? 3 - z/4 = 8 A. z = –44 B. z = –20 C. z = 20 D. z = 44

OpenStudy (ajspeller):

Multiply each term by 4 to remove the fraction and then solve for z.

OpenStudy (anonymous):

C subtract 3 from each side then multiply each side by 4

OpenStudy (anonymous):

\[\frac{ 3-z }{ 4 }=8\] Was this the original equation?

OpenStudy (anonymous):

or this:\[3-\frac{ z }{ 4 }=8\]

OpenStudy (anonymous):

No

OpenStudy (anonymous):

The 2nd one

OpenStudy (anonymous):

In that case the answer is z=-20 Multiply each term by 4 \[4(3)-4(\frac{ z }{ 4 })=4(8)\] You get \[12-z=32\] Then add z to each side and subtract 32 from each side and you get z=-20